OpenStudy (anonymous):

how can you solve the equation 4^x=2^x+2 graphically?

jimthompson5910 (jim_thompson5910):

graph the two equations y = 4^x y = 2^x+2 on the same coordinate axis

jimthompson5910 (jim_thompson5910):

where they cross, or intersect, is the point that leads to the solution (the x coordinate of that point is the solution to the original equation)

OpenStudy (mathmale):

K: Have you graphed exponential functions before? Both of these are exponential functions. For y = 4^x we could make a table of values: x y=4^x Points --- --------- ------- 0 1 (0,1) 1 4 2 16 Could you graph these points?